### v1

# The reason peculiar velocities grow faster in general relativity than in Newtonian gravity

Abstract
An increasing number of surveys has been reporting large-scale peculiar motions with sizes and speeds in excess of those allowed by the concordance cosmological model. These are the so called bulk flows, the presence of which has come to be treated as a problem for the CDM paradigm. However, the limits of the CDM model are based on Newtonian studies, which predict the mediocre growth-rate for the peculiar-velocity field (). Recently, a few fully relativistic treatments have appeared in the literature, arguing for a much stronger velocity growth that could explain the reported fast and deep bulk flows. What separates the Newtonian from the relativistic studies is the gravitational input of the peculiar flux, namely of the kinetic energy triggered by the moving matter. The latter has no direct gravitational contribution in Newtonian theory, but it does so in general relativity. This drastically changes the driving agent of the peculiar-velocity field and boosts its linear growth. The aim of this work is to directly compare the two treatments, as well as identify and discuss the reasons for their different results. In the process, we also demonstrate how one could recover the relativistic growth-rate from a Newtonian setup by selectively including certain (typically ignored) source-free terms into the Poisson equation. This way, we provide a unified covariant comparison of the Newtonian, the quasi-Newtonian and the fully relativistic studies.
1 Introduction
Large-scale peculiar motions, commonly known as bulk flows, are an observational certainty, confirmed by numerous surveys over several decades. These surveys largely agree on the direction of the reported bulk flows, but they seem to disagree on their size and speed. Moreover, while many of the reported bulk flows have sizes and velocities within the theoretical boundaries of the current cosmological model (e.g. see [1]-[4], there has been an increasing number of recent surveys claiming bulk peculiar flows considerably faster and deeper than it is theoretically expected (e.g. see [5]-[9]). To these, one could also add the controversial “dark flows”, with sizes and velocities well in excess of the CDM limits [10, 11]. Typically, the surveys that agree with the CDM expectations cover scales smaller than 100 Mpc, whereas those that disagree extend beyond the aforementioned threshold. As yet, the reason responsible for this discrepancy remains unknown and a subject of debate within the cosmological community. After all, if confirmed, the bulk-flow question could prove a serious problem for the concordance CDM model. The reader is referred to [12] for a recent review on cosmological peculiar motions, with more details and references.
When it comes to peculiar motions, the CDM predictions have been built upon Newtonian theory. The latter argues for a relatively mild linear growth-rate of for the peculiar-velocity field () [13, 14]. This was also the result of the so-called “quasi-Newtonian” treatments [15, 16], which however reduce to mere Newtonian despite their initially relativistic profile. One of the reasons for remaining within the Newtonian limits was probably the anticipated technical difficulty of addressing the bulk-flow question relativistically. Another was the widespread expectation that, when dealing with non-relativistic peculiar motions, Einstein’s theory was unlikely to drastically change the existing Newtonian picture. It is therefore not surprising that fully relativistic linear studies of peculiar velocities have only recently appeared in the literature [17]-[19]. These treatments are arguing for a considerably stronger peculiar growth-rate, namely for at a minimum, which has the potential of addressing both the higher speeds and the greater depths of the bulk flows reported in [5]-[9] for example. The reason for the profound difference between the results of the Newtonian/quasi-Newtonian and the relativistic studies can be tracked down to the fundamentally different way the two theories treat both the gravitational field and its sources. More specifically, to the entirely different role played by the “peculiar flux” in these two types of theoretical approach.
Peculiar flows are nothing else but matter in motion and moving matter means nonzero energy flux (e.g. see [20, 21]). In general relativity, as opposed to Newtonian gravity, energy fluxes contribute to the stress-energy tensor and therefore gravitate [17]. The gravitational input of the peculiar flux, which survives even at the linear perturbative level, is not accounted for neither in the Newtonian treatments nor in their quasi-Newtonian counterparts, though for different reasons. In the former studies this happens by default, while in the latter due to the strict constraints the quasi-Newtonian approximation imposes upon the perturbed spacetime. These compromise the relativistic nature of the approach and reduce it to a mere Newtonian for all practical purposes. Although the problematic nature of the quasi-Newtonian approach has been known (e.g. see related “warning” comments in § 6.8.2 of [21]), this has not been widely appreciated. The likely reason is the absence of a direct comparison with a fully relativistic analysis (until very recently - see [22] for such a comparison).
In a fully relativistic linear study, the gravitational input of the peculiar flux feeds into the associated conservation laws and subsequently into the linear evolution formulae of the peculiar-velocity field. In so doing, the relativistic treatment brings into play the effects of density perturbations, namely of structure formation, in addition to those of the peculiar flux. All this drastically modifies the agent driving the linear peculiar velocities. As a result, unlike its Newtonian/quasi-Newtonian counterpart, the relativistic driving agent no longer decays with time. This in turn ensures that the relativistic differential equations and their associated solutions differ considerably. More specifically, the general relativistic analysis supports a substantially stronger linear growth for peculiar velocities during the Einstein-de Sitter phase of the universe. We demonstrate here that, even in the “minimalist” scenario where only the gravitational input of the peculiar flux is taken into account, the velocity growth-rate is . Moreover, we also show that the latter can readily increase to the -rate of [17]-[19] when the effects of structure-formation are also accounted for. In either case, the resulting velocity field has the potential of explaining the speed and the depth of the fast bulk flows reported in [5]-[9]. Put another way, general relativity could naturally relax the typical (Newtonian) CDM limits to accommodate the aforementioned bulk peculiar flows.
This work provides a direct comparison between the Newtonian/quasi-Newtonian and the relativistic analysis of linear peculiar velocities on a Friedmann-Robertson-Walker (FRW) background with flat spatial sections. We set the theoretical framework of the two approaches, identify the point where they start to diverge from each other and explain why this happens. We then derive the linear evolution laws of the associated peculiar-velocity fields, highlight the differences between the corresponding solutions and discuss their implications for addressing the ongoing bulk-flow question. In the process, we also highlight the role of the acceleration/4-acceleration as the driving force of linear peculiar velocities. More specifically, we show that the stronger the 4-acceleration, the faster the driven peculiar-velocity field. Finally, we demonstrate a way of recovering the relativistic results from the Newtonian setup, by selectively accounting for certain source-free terms in Poisson’s formula that are typically sidestepped. Although, this recovery is rather ad hoc and not uniquely determined it provides useful analogies between the two types of study. For instance, one could argue that ignoring the above mentioned source-free terms in the Newtonian studies, is analogous to bypassing the purely general relativistic gravitational input of the peculiar flux.
We start our presentation with the traditional study of cosmological linear peculiar velocities, which employs Newtonian gravity and proceeds along the lines of [13, 14] and [23] in section § 2. We provide the covariant analogue of the Newtonian analysis in § 3.3.1, in order to familiarise the reader with the covariant formalism and connect the standard Newtonian treatment of § 2 with the fully relativistic approach of § 3.3.3. There, we also discuss the physical reasons the Newtonian/quasi-Newtonian studies of linear peculiar velocities arrive at considerably weaker results, compared to those of the relativistic treatment. In so doing, we highlight the key role played by the 4-acceleration, keeping in mind that it is the driving force of the linear peculiar-velocity field (see § 3.3.4). We also outline a way of bridging the aforementioned differences between the Newtonian and the relativistic studies, by (selectively) including extra contributions to the Newtonian gravitational potential (see § 4). Finally, we close with the discussion of our results and of their role in the quest for an answer to the bulk-flow puzzle.
2 Newtonian treatment of peculiar velocities
Let us employ Newtonian theory and adopt comoving coordinates . These are related to their corresponding physical coordinates by means of , where is the cosmological scale factor. Then, the velocity () of a typical observer in the Newtonian version of a perturbed Einstein-de Sitter universe decomposes as
with representing the Hubble parameter. Also, is the observer’s peculiar velocity relative to the universal rest-frame, which has been typically identified with the rest-frame of the Cosmic Microwave Background (CMB). Within the adopted cosmological framework, the linear evolution of the peculiar-velocity field is monitored by the associated Euler equation, namely by [13]
where is the peculiar gravitational potential and is the gradient in comoving coordinates.333For the comoving analogue of expression (2), the interested reader is referred to Eq. (5a) in [23]. The latter acts as a linear source of peculiar-velocity perturbations and satisfies Poisson’s formula, that is
with being the familiar density contrast, the matter density and its background value.
Differentiating Eq. (2) in time, recalling that and to zero order, while keeping up to linear order terms, leads to
One may ignore the right-hand side of the above, by assuming that the gradient of the peculiar gravitational potential varies slowly in time for example. Then, (4) reduces to
which accepts the power-law solution
on all scales. Therefore, according to Newtonian physics, linear peculiar velocities are expected to grow as after recombination and throughout the Einstein-de Sitter epoch.
One can also refine solution (6) by incorporating to it the source term on the right-hand side of (4). Following [13, 14], we start by recalling that the peculiar gravitational acceleration is
Given that and also depend on time, after a relatively straightforward calculation, the temporal derivative of the above leads to the linear auxiliary relation
Note that in deriving this relation we have used the background expressions and , together with the linear result . The latter follows from the continuity equation, which reads at the linear level, with and being the perturbed matter density and Hubble parameter respectively. Setting and , substituting into the linear continuity equation and keeping up to first-order terms, we obtain the desired relation, namely at . Finally, combining (2), (4) and (8), we arrive at the differential equation
and to the associated power-law solution [13, 14]
on all scales. Comparing solutions (6) and (10) makes it clear that the inclusion of the nonzero right-hand side of (4) has no physical/practical significance. The rate of the growing mode in the former of the two solutions has remained the same and only the decaying mode has changed.
Our last note is on the peculiar acceleration ( - see Eq. (7) above), which acts as the driving force of the linear -field in Newtonian theory. Combining (2) with solution (10), one immediately realises that the Newtonian driving force decays with time as
since after matter–radiation equality. In section § 3.3.4, we will demonstrate that the same time-decay law also applies to the 4-acceleration that drives linear peculiar velocities in the quasi-Newtonian studies. In contrast, the 4-acceleration driving the peculiar-velocity field in the relativistic treatments is either constant (at a minimum), or increases with time (see § 3.3.4 below). This difference offers another way of explaining why the Newtonian/quasi-Newtonian studies lead to the mediocre growth-rate and why general relativity supports faster and deeper bulk flows that it has been largely anticipated.
3 Covariant treatment of peculiar velocities
The reported large-scale peculiar motions are believed to have started as weak peculiar-velocity perturbations that were amplified to the observed bulk flows by structure formation. In this section we will provide a “unified” covariant study looking into the linear evolution of peculiar velocities in cosmology. In so doing, we will also facilitate the direct comparison between the Newtonian, the quasi-Newtonian and the relativistic treatments of the subject.
3.1 CMB and matter frames
Studies of peculiar velocity require two relatively moving coordinate systems, with one of them following the peculiar motion of the matter and the other the reference frame of the universe. The latter has been typically identified with the coordinate system of the CMB, namely the frame relative to which we define and measure peculiar motions.
In Newtonian theory, the velocities of the aforementioned pair of observers are related via the familiar Galilean transformation, namely by
where and are the velocities of bulk-flow observers and of their CMB partners respectively. Also, the peculiar velocity of the former with respect to the latter (with ). In relativity, the Galilean transformation is replaced by the (also familiar) Lorentz boost, which in the case of non-relativistic peculiar motions reads
Here, in contrast to (12), and are the (timelike) 4-velocities of the aforementioned observers, while is the (spacelike) velocity of the peculiar motion (with ). Put another way, and by construction. Also, when dealing with non-relativistic peculiar flows. Note that, although the transformations (12) and (13) may seem formally identical, they are different because the former involves spacelike vectors only, whereas in the latter the two 4-velocity vectors are timelike.
3.2 The key role of the peculiar flux
The key difference between the Newtonian and the relativistic study of peculiar motions follows from the fact that, in relativity, the nature of the matter fields changes between the relatively moving frames/observers. In particular, a cosmic medium that appears perfect to one reference frame, will appear imperfect to any other relatively moving coordinate system. The imperfection takes the form of an effective energy flux vector, solely triggered by relative-motion effects. More specifically, let us consider a perturbed Friedmann universe, with zero spatial curvature. Then, setting the pressure (both the isotropic and the viscous) to zero, we arrive at the linear relations
and
between the CMB and the bulk-flow (i.e. the tilded) frames. In the above , , and are respectively the density the isotropic pressure, the viscosity and the energy flux of the matter relative to the CMB frame, with their tilded analogues measured in the matter frame. Following (14), the density and the pressure (both the isotropic and the viscous) of the matter are the same in the two coordinate systems at the linear level. On the other hand, according to (15), there is a difference between the energy fluxes entirely due to relative-motion effects. In fact, Eq. (15)) ensures that, even if the flux vanishes in one frame, there is a nonzero peculiar flux in any other relatively moving coordinate system, entirely due to relative-motion effects. Clearly, one cannot set the flux to zero in both frames simultaneously without switching the peculiar velocity off as well. Put another way, in the presence of peculiar motions, there is always a nonzero energy flux in the system, which in turn means that the cosmic medium is no longer perfect.
Assuming vanishing flux in the matter frame, we may set . Then, Eq. (15) guarantees a nonzero peculiar flux equal to in the CMB frame. The latter survives at the linear perturbative level and, in relativity, it contributes to the perturbed energy-momentum tensor and therefore to the relativistic gravitational field. As a result, the linear conservation laws of the energy and the momentum densities acquire flux-related terms and take the forms
respectively. In the above, is the 3-divergence of the -field (with in the background), is the associated 4-acceleration and is the 3-D covariant derivative operator. Note that, according to (16b), the 4-acceleration is nonzero, despite the absence of pressure. This is a direct consequence of the fact that, in the presence of relative motions, there is always a nonzero energy flux in the system and the cosmic medium can no longer be treated as perfect.444The linear expression (15) guarantees that nonzero peculiar flux ensures a nonzero 4-acceleration as well. The only exception is when , which however corresponds to a “set of measure zero” in probabilistic terms. Moreover, this one and only exception leads to a -field that decays as . Since peculiar velocities start weak at recombination, there should be no bulk flows today if they were to decay in time. Recall that, even in Newtonian theory, linear peculiar velocities grow as – see [13, 14]. As it turns out (see § 3.3 next), it is the peculiar-flux input to the gravitational field that separates the relativistic studies of peculiar motions from the rest.
3.3 Newtonian/quasi-Newtonian vs relativistic treatments
As mentioned above, the observed bulk peculiar flows begun as peculiar velocity perturbations, which were subsequently amplified by structure formation. The driving agent of the amplification process depends on the theory that one employs to study the evolution of the peculiar-velocity field. In Newtonian physics, the driving force is gravity and more specifically the gravitational acceleration. Mathematically, this is reflected in the linear evolution equation of the -field, which reads
with being the Newtonian gravitational potential. In relativity, gravity is not a force, but the manifestation of spacetime curvature. As a result, in a relativistic treatment, linear peculiar velocities are driven by non-gravitational forces and the evolution of linear peculiar velocities is monitored by
where is the 4-acceleration measured in the reference CMB frame.555It is straightforward to show that expression (18) also follows from (16b) after substituting in the right-hand side of the latter. Also note that in the Newtonian equations overdots imply convective derivatives (e.g. ), while in their relativistic counterparts overdots indicate covariant differentiation along the associated 4-velocity field (e.g. ). It is therefore clear that the linear evolution of the -field is determined by the form and the impact of the associated driving agent.
3.3.1 Newtonian treatment
Newtonian peculiar-velocity studies start from differential equation (17). Recalling that to zero order, the linearised time derivative of the latter leads to
since to first approximation. As noted in § 2 before, ignoring the right-hand side of the above, leads to the power-law solution . However, by appealing to the Jeans’ swindle (e.g. see [24]) and writing Poisson’s equation as ,666Hereafter, the gravitational constant , as well as the velocity of light, have been set to equal unity. a relatively lengthy but fairly straightforward calculation yields
which substituted into the right-hand side of (19) gives
with in the background. The above solves to give
on all scales and in complete agreement with solution (10) in § 2 and also with [13, 14]. Therefore, as expected, incorporating the right-hand side of (19) into the calculation has only changed the decaying mode of the solution.
3.3.2 Quasi-Newtonian treatment
Expression (18) is also the starting point of the quasi-Newtonian studies of linear peculiar velocities in cosmology [15, 16]. However, the agreement with the relativistic analysis stops at this point because of the strict constraints imposed upon the perturbed spacetime by the quasi-Newtonian approximation. These constraints compromise its relativistic nature of the analysis and eventually lead to Newtonian equations and results (see § 6.8.2 in [21] for related “warning” comments). More specifically, the quasi-Newtonian approach starts by setting the vorticity to zero at the linear level. This has the benefit of simplifying the mathematics, because it allows one to express the 4-acceleration as the spatial gradient of a scalar potential (), so that
There is a downside, however, because is an ad hoc potential, with no expression to describe it, while its linear evolution follows from the (non-uniquely determined) ansatz [15]. In addition, the aforementioned ansatz requires setting the linear shear to zero as well. There is more to it however. When it comes to the study of peculiar velocities, the most serious downside of the quasi-Newtonian treatment is that it inadvertently bypasses the gravitational contribution of the peculiar flux. This severely compromises the relativistic nature of the study and makes it identical to its purely Newtonian counterpart for all practical purposes. Indeed, combining (18) and (23) the differential equation governing the linear evolution of the peculiar-velocity field takes the form [15, 16]
and accepts the power-law solution
both of which are identical to their purely Newtonian analogues (compare to Eqs. (21) and (22) in § 3.3.1 earlier). Therefore, according to the Newtonian and the quasi-Newtonian studies, linear peculiar velocities grow no faster than during the Einstein-de Sitter era. However, this growth-rate is not strong enough to explain the fast and deep bulk flows reported by several surveys over the last fifteen years or so.
3.3.3 Relativistic treatment
The relativistic analysis of linear peculiar velocities imposes no constraints on the perturbed spacetime and, crucially, it also accounts for the gravitational input of the peculiar flux [17, 18, 19]. The latter feeds into Einstein’s equations and emerges into the equations monitoring a cosmological model endowed with a peculiar-velocity field. As a result, there is no longer need neither to introduce an ad hoc scalar potential () for the 4-acceleration (see Eq. (23) above), nor adopt the ansatz for its time evolution. Instead, the 4-acceleration emerges naturally in the linear formulae governing the evolution of the density and the expansion gradients. These are respectively given by [22]
and
with and representing perturbations in the matter density and the universal expansion respectively.777Expressions (26) and (27) follow by respectively linearising the spatial gradient of the energy conservation law (see (16) in § 3.2) and that of the Raychaudhuri equation [22]. Alternatively, one can obtain (26) and (27) by linearising the nonlinear formulae (2.3.1) and (2.3.2) of [20], or Eqs. (10.101) and (10.102) of [21], while taking into account the linear relation (16b) and that in our case . Also, is the spatial divergence of the peculiar velocity. The above is the relativistic set of differential equations that governs the linear evolution of peculiar velocities in a perturbed Einstein-de Sitter universe. It is therefore clear that the relativistic analysis naturally couples the driving force of the peculiar-velocity field, namely the 4-acceleration, to both the density and the expansion gradients.
Differentiating (26) with respect to time, substituting the right-hand side of (27) in the resulting expression and then using Eq. (18), we arrive at [22]
which is the relativistic formula governing the evolution of linear peculiar velocities in an Einstein-de Sitter cosmology. Note that in deriving the above we have also used the background relation , together with the linear commutation laws and . Expression (28) is a non-homogeneous differential equation, the homogeneous component of which reads888The terms homogeneous/non-homogeneous refer to the nature of the differential equation and not to the homogeneity, or not, of the host spacetime. The latter is both inhomogeneous and anisotropic at the linear level.
and accepts the solution
on all scales [22]. Therefore the relativistic analysis has led to the considerably stronger growth-rate of for linear peculiar velocities. Moreover, according to the theory of differential equations, solution (30) provides the minimum growth rate of the -field. Indeed, a fairly well known theorem states that the full solution of a non-homogeneous differential equation forms from the general solution of its homogeneous part (in our case from (30)) and from a partial solution of the full equation. Therefore, solving (28) in full will make physical difference only if its partial solution grows faster than the fastest growing mode of the homogeneous solution. Put another way, the power-law is the minimum growth-rate of the linear peculiar-velocity field. Indeed, by following an approach that accounted for some (though not for all) of the effects of the density gradients seen on the right-hand side of (28), the studies of [17]-[19] have led to the growth-rate of for the linear -field. Overall, the relativistic analysis has led to considerably stronger growth-rates for peculiar velocities than the Newtonian/quasi-Newtonian treatments. This means that general relativity supports faster and deeper residual bulk flows than it is generally anticipated. Perhaps bulk flows like those reported in the surveys of [5]-[9] for example.
3.3.4 On the role of the 4-acceleration
A complementary, less technical, but perhaps more physically intuitive way of demonstrating and explaining the faster growth-rates of the relativistic analysis, is by looking closer at the role of the peculiar 4-acceleration. According to the linear theory, the latter is the driving force of the -field (see Eq. (18) in § 3.3 earlier). Therefore, the stronger the 4-acceleration, the faster the peculiar velocity.
Following Eqs. (26) and (27), the input of the peculiar-flux to the relativistic gravitational field also couples the peculiar 4-acceleration to the density and the expansion gradients. Let us ignore for the moment the effects of the aforementioned gradients and keep only that of the peculiar flux. Then, on using (18), expressions (26) and (27) reduce to
which combine to the linear relation
Taking the time derivative of (31a), employing Eq. (32) above and using the auxiliary background and linear relations given in § 3.3.3 previously, one can easily show that . Accordingly, when only the flux-effects are accounted for, linear peculiar velocities are driven by a time-invariant 4-acceleration. In such case, the differential equation (18) accepts the linearly growing solution
on all scales and in full agreement with solution (30) obtained in [22] and in § 3.3.3 above.
Keeping the density and the expansion gradients on the right-hand side of (26) and taking its time derivative, we arrive at the following propagation formula for the linear peculiar 4-acceleration [17]
This is a non-homogeneous differential equation, the homogeneous part of which solves to give (recall that in the background). Note that, in line with the theorem on differential equations mentioned earlier, the mode provides the minimum growth-rate of the peculiar 4-acceleration. Finally, after substituting the latter result into the right-hand side of Eq. (18), we find that
at a minimum and in full agreement with [17] (see also [18, 19]). Consequently, when some of the gradient effects are also accounted for, the driving force of the peculiar-velocity field is no longer constant, but grows in time (as ). As a result, the associated -field grows faster than linearly, namely as . It is also conceivable that, if all the structure-formation effects were to be accounted for, the peculiar-velocity field may grow even faster.
In contrast to its relativistic counterparts, the quasi-Newtonian 4-acceleration decays in time as [15]. On using (18), this leads to a linear peculiar-velocity field that grows as only. It goes without saying that the purely Newtonian studies lead to the same results exactly (e.g. see Eq. (11) in § 2). The fact that, in the Newtonian and quasi-Newtonian studies, the linear peculiar velocities are driven by a decaying force, whereas in the relativistic treatments the driving agent is constant, or even increasing in time, explains both the rather mediocre growth-rate of the former treatments and the substantially stronger (, or even ) rates of the latter.
3.3.5 On the peculiar vorticity and shear
A generic peculiar-velocity field is expected to have nonzero vorticity and shear. The latter are respectively defined as the antisymmetric and the symmetric traceless components of the -field. In other words, is the peculiar vorticity and is the peculiar shear. These kinematic variables do not necessarily evolve as the peculiar velocity itself. Indeed, taking the spatial gradient of (29) and using the linear commutation laws and , we arrive at the following differential equation
for the linear evolution of the peculiar vorticity. Clearly, an exactly analogous differential formula monitors the linear peculiar shear (). Recalling that during the Einstein-de Sitter epoch, the above solves to give
with the same solution governing the evolution of the peculiar shear () as well. Therefore, both and grow considerably slower than the peculiar-velocity field itself (recall that – see solution (30) in § 3.3.3 before). Note that this result is in agreement with the one obtained by the earlier relativistic analysis of [17].
3.3.6 On the impact of the late-time acceleration
So far our analysis has been confined to the Einstein-de Sitter epoch of the universe. Within the framework of the CDM scenario the earlier dust-dominated era is followed by a recent phase of accelerated expansion, where the energy density of the cosmos is typically dominated by a cosmological constant, or by dark energy. Therefore, our results apply to the period between matter–radiation equality and the onset of cosmic acceleration. Once the latter epoch has started, the evolution of the peculiar-velocity field is expected to change drastically. More specifically, the accelerated expansion is expected to suppress the growth of peculiar velocities, in the same way it suppresses essentially all types of perturbations, like those in the matter density for example. This claim is also supported by numerical studies showing that dark energy inhibits the growth of peculiar velocities compared to the Einstein-de Sitter epoch [25]. Moreover, complementary studies looking into the linear evolution of peculiar velocities during a phase of de Sitter inflation, showed exponential decay for the -field [18]. Nevertheless, some at least of the strong earlier growth of the Einstein-de Sitter period should survive through the accelerated phase to the present. On these grounds, one should expect to observe residual bulk flows faster and deeper than those anticipated by the Newtonian studies, but also to see a drop in their mean velocity at lower redshifts [22]. Intriguingly, such a late-time suppression of the -field has been reported in the surveys of [8, 9].
4 An effective gravitational potential
As we have already mentioned, it is possible to recover the relativistic solution through the Newtonian analysis, although the mapping is neither unique nor physically equivalent in a strict sense. In so doing, one needs to introduce a specific ansatz for the evolution of the gravitational potential (or, equivalently, for the integration constants that appear when Poisson’s equation is used in reverse). More specifically, in line with the covariant version of the Newtonian study (see § 3.3.1 earlier), the divergence of Eq. (20) gives
which, upon integrating once in space, yields
Here could (in principle) have both temporal and spatial dependence and encodes the solenoidal (divergence-free) integration mode that is left undetermined by the divergence in Eq. (38) and is fixed only after specifying boundary conditions and/or a coordinate choice. Substituting the above into the right-hand side of (19), one arrives at the following (non-homogeneous) differential equation
for the linear evolution of the peculiar-velocity field. Expression (40) differs from Eq. (21) only due to its nonzero right-hand side, the presence of which allows one to change the standard Newtonian evolution of the linear peculiar velocities: allowing , we effectively parametrize the freedom in the time evolution of not fixed by Poisson alone. For instance, choosing , one can recover the relativistic growth-rate of obtained in § 3.3.3 before (see Eqs. (29) and (30) there). It goes without saying, however, that our choice of is not uniquely determined and therefore this procedure should be viewed as an illustrative mapping rather than as a physical equivalence between Newtonian gravity and general relativity.
It is worth pointing out that a vector function analogous to was obtained in [14], also after involving and integrating Poisson’s formula. There, the aforementioned vector was set to zero by an appropriate choice of coordinates or boundary conditions. This is reminiscent of the fact that the relativistic peculiar flux is also a vector that can be set to zero (albeit only locally) by moving to the energy (or Landau-Lifshitz) frame. Moreover, even in the energy frame, the peculiar energy flux does not disappear, but re-emerges “disguised” as “peculiar particle flux”. We remind the reader that, in relativity and depending on the specifics of the problem in hand, one may adopt the energy frame, where the energy flux vanishes but there is nonzero particle flux. Alternatively, one may choose to use the particle (or Eckart) frame, where the situation is reversed (e.g. see [20, 21]).
5 Discussion
Large-scale peculiar motions, commonly referred to as bulk flows, are commonplace in our universe, having been repeatedly verified by many surveys (e.g. see [1]-[9] for a representative, though incomplete, list). Although currently the aforementioned flows typically encompass domains of up to several hundred Mpc, we believe that they have started as weak peculiar-velocity perturbations around the time of recombination, which grew larger and faster by structure formation and by the increasing non-uniformity of the post-recombination universe. Be that as it may, the evolution and the implications of these bulk peculiar motions are still largely unknown. Moreover, during the last fifteen years or so, the bulk-flow puzzle has become even more perplexing, after an increasing number of surveys have reported sizes and speeds well in excess of those allowed by the concordance cosmological model, namely by the CDM paradigm. In fact, if one is allowed to take into account the controversial dark flows, with sizes of the order of 1000 Mpc and speeds close to 1000 km/sec (see [10, 11] for example), the situation becomes rather untenable.
Having said that, the CDM constraints on the speed and the depth of the observed bulk peculiar flows are based entirely on Newtonian theory. There is a serious difference, however, between Newtonian gravity and general relativity when it comes to the study of peculiar motions, which stems from the fundamentally different way the two theories treat both the gravitational field and its sources. To begin with, recall that bulk flows are nothing else but matter in motion and that moving matter means nonzero energy flux. In Newton’s theory only the density of the matter contributes to the gravitational field, via Poisson’s equation, while any energy fluxes that may exist have no gravitational input. In relativity, on the other hand, the pressure (both isotropic and viscous), as well as the fluxes also contribute to the energy-momentum tensor and therefore to the (relativistic) gravitational field. It is the input of the peculiar flux that has been bypassed in essentially all the mainstream studies so far. So, accounting for the gravitational contribution, or not, is the key difference that separates the relativistic from the Newtonian (as well as the quasi-Newtonian) treatments of cosmological peculiar motions.
Without incorporating the gravitational input of the peculiar flux, the Newtonian and quasi-Newtonian analyses have both led to the same mediocre () growth-rate for the linear peculiar-velocity field [13]-[16]. Unfortunately, the latter is too slow to explain the reported fast bulk flows. On the other hand, by including the peculiar-flux contribution to the gravitational field alone, the relativistic analysis has led to the considerably faster growth-rate of (see § 3.3.3 here and also [22]). Moreover, the standard theory of differential equations guarantees that is the minimum growth supported by the relativistic studies. To a large extent, the latter was already foreseen in the studies of [17]-[19], which arrived at the faster linear growth-rate of , after accounting for some of the inhomogeneity effects as well.
An alternative and also complementary explanation of the aforementioned relativistic results involves the acceleration/4-acceleration, which are the driving forces of the linear peculiar-velocity field in all three approaches. Then, a closer inspection (see § 3.3.4 above) revealed that the driving agent decays as in both the Newtonian and the quasi-Newtonian studies. In contrast, the relativistic 4-acceleration is either constant in time (at a minimum - see [22] and § 3.3.3 here), or it was found to increase as throughout the Einstein-de Sitter phase of the universe (see [17]-[19]). This drastic difference explains why general relativity supports significantly faster peculiar-velocity fields. Also note that, within the framework of the standard CDM model, cosmic acceleration starts late into the Einstein-de Sitter phase of the universe. Then, some (at least) of the increased velocity growth taking place during the Einstein-de Sitter epoch should survive through the subsequent accelerated phase and, in so doing, lead to faster and deeper residual bulk flows today. Perhaps like those reported in [5]-[9] for example. Moreover, all this happens within the standard CDM model and without the need for any new free parameters.
The exact analytical results of the linear analysis presented here, as well as in the earlier works of [17]-[19] and [22], can also provide a useful testing ground for the emerging numerical studies simulating the post-recombination universe. The fact that all the aforementioned theoretical treatments are gauge-invariant is an additional advantage, since their results are free of any gauge-related ambiguities.999The peculiar-velocity field () vanishes by default in the unperturbed FRW background, which makes it a gauge-invariant linear perturbation [26]. Recall that (to the best of our knowledge) the numerical simulations rely on particular gauge choices when addressing the problems in hand.
As yet, most of the simulations that involve relativistic techniques do not consider the evolution of peculiar velocities, with a couple of exceptions perhaps that looked at the effects of the peculiar vorticity (e.g. see [27]). The latter was found to have a negligible impact, which was to be expected since it grows considerably slower that the peculiar velocity itself (even in the relativistic analysis – see [17] and also § 3.3.5 here). Moreover, the simulations typically operate within the CDM paradigm (e.g. see [25] and also [28] for a recent review). So, the numerical codes are adapted to the recent accelerated phase of the universe where the peculiar-velocity field is suppressed, instead of the preceding Einstein-de Sitter epoch where peculiar velocities are strongly enhanced. As a result, so far, there is no common ground to facilitate the direct comparison between the analytical and the numerical work. Nevertheless, the relativistic simulations are still under development and there is room for testing and improvement, especially in view of repeated surveys reporting bulk flows in excess of the (Newtonian-based) CDM limits. Hopefully, the results of the analytical relativistic studies will provide the motivation for extending and adapting the numerical codes to address the ongoing bulk-flow puzzle.
In summary, general relativity can provide a simple and physically motivated theoretical solution to the ongoing bulk-flow puzzle, by just taking into account the (purely general relativistic) gravitational contribution of the peculiar flux. This can in turn relax the current (purely Newtonian) peculiar-velocity expectations of the CDM model and, in so doing, allow the latter to naturally accommodate bulk flows faster and deeper than anticipated, like those reported in [5]-[9] for example. Then, the perceived discrepancy between the current cosmological model and the aforementioned bulk-flow reports may not reflect a generic problem in the CDM paradigm, but instead it may indicate the use of the inappropriate gravitational theory in our studies of cosmological peculiar flows.
Acknowledgements: CGT acknowledges support from the Hellenic Foundation for Research and Innovation (H.F.R.I.), under the “First Call for H.F.R.I. Research Projects to support Faculty members and Researchers and the procurement of high-cost research equipment Grant” (Project Number: 789).

### v2

@@
 This work provides a direct comparison between the Newtonian/quasi-Newtonian and the relativistic analysis of linear peculiar velocities on a Friedmann-Robertson-Walker (FRW) background with flat spatial sections.
-We set the theoretical framework of the two approaches, identify the point where they start to diverge from each other and explain why this happens.
+The linear nature of all the studies implies that their results formally apply to relatively large scales, typically near and beyond the 100 Mpc threshold, where one feels confident enough to ignore the nonlinear effects.
+We begin by setting the theoretical framework of the approaches, identify the point where they start to diverge from each other and explain why this happens.
 We then derive the linear evolution laws of the associated peculiar-velocity fields, highlight the differences between the corresponding solutions and discuss their implications for addressing the ongoing bulk-flow question.
@@
 One can also refine solution (6) by incorporating to it the source term on the right-hand side of (4). Following [13, 14], we start by recalling that the peculiar gravitational acceleration is
-Given that and also depend on time, after a relatively straightforward calculation, the temporal derivative of the above leads to the linear auxiliary relation
+Given that and also depend on time, after a relatively straightforward calculation, the temporal derivative of the above leads to the linear time-evolution equation for , namely
 Note that in deriving this relation we have used the background expressions and , together with the linear result .
@@
 3.3.1 Newtonian treatment
-Newtonian peculiar-velocity studies start from differential equation (17). Recalling that to zero order, the linearised time derivative of the latter leads to
+Newtonian peculiar-velocity studies start from differential equation (17). Recalling that to zero order, the linearised time derivative of the former leads to
 since to first approximation.
@@
 More specifically, the accelerated expansion is expected to suppress the growth of peculiar velocities, in the same way it suppresses essentially all types of perturbations, like those in the matter density for example.
-This claim is also supported by numerical studies showing that dark energy inhibits the growth of peculiar velocities compared to the Einstein-de Sitter epoch [25].
-Moreover, complementary studies looking into the linear evolution of peculiar velocities during a phase of de Sitter inflation, showed exponential decay for the -field [18].
-Nevertheless, some at least of the strong earlier growth of the Einstein-de Sitter period should survive through the accelerated phase to the present.
-On these grounds, one should expect to observe residual bulk flows faster and deeper than those anticipated by the Newtonian studies, but also to see a drop in their mean velocity at lower redshifts [22].
-Intriguingly, such a late-time suppression of the -field has been reported in the surveys of [8, 9].
+Analytical relativistic studies that directly address and quantify the impact of a late CDM phase on large-scale peculiar velocities would be very useful indeed.
+At present, there is support for the aforementioned suppressing effect of the universal acceleration, but indirect.
+In particular, according to the Newtonian numerical work of [25], dark energy inhibits the growth of peculiar velocities, when compared to the Einstein-de Sitter epoch.
+Also, studies looking into the linear evolution of peculiar velocities during a phase of de Sitter inflation, showed exponential decay for the -field [18].
+Nevertheless, even if peculiar velocities are suppressed during the late -dominated phase, some at least of the strong growth of the preceding Einstein-de Sitter period should survive through to the present.
+On these grounds, one should expect to measure bulk velocities faster than anticipated by the Newtonian studies at higher redshifts, but also to see a decrease in their mean value at lower redshifts (see [22] for further discussion).
+Here, we note that a peculiar-velocity profile, with increased values at higher redshifts (i.e.
+at earlier times) and a late-time suppression has been reported in the survey of [6] and (more pronounced) in those of [8, 26].
+Put another way, our work appears to provide theoretical support to the aforementioned observational studies.
 4 An effective gravitational potential
@@
 Hopefully, the results of the analytical relativistic studies will provide the motivation for extending and adapting the numerical codes to address the ongoing bulk-flow puzzle.
+Observationally, the time evolution of the peculiar velocities can be mapped into a redshift evolution at low , enabling a comparison between our theoretical predictions and the standard Newtonian expectation using peculiar-velocity surveys such as CosmicFlows-4.
+However, peculiar-velocity measurements are subject to a well-known degeneracy between the assumed background cosmology and local kinematic contributions.
+This degeneracy can be partially mitigated through independent distance indicators, such as Type Ia supernovae or Tully-Fisher relations, which provide redshift-independent distance estimates.
+Nevertheless, the use of such indicators introduces additional uncertainties that grow with redshift, thereby impacting the reconstruction of the peculiar-velocity field and limiting the robustness of the physical inferences.
+A detailed implementation of unbiased and statistically robust estimators — such as the minimum-variance reconstruction techniques applied to CosmicFlows-4 [8, 26] — is beyond the scope of the present work and will be addressed in future studies.
 In summary, general relativity can provide a simple and physically motivated theoretical solution to the ongoing bulk-flow puzzle, by just taking into account the (purely general relativistic) gravitational contribution of the peculiar flux.